OpenStudy (pedro):

Emma finished a small bag of her favorite snack that weighed 0.6 pound. Wendy ate 2/7 of a one-pound bag of the snack. Who ate more? Explain.

OpenStudy (turingtest):

0.6=6/10=42/70 2/7=20/70 So it looks like Emma ate more

OpenStudy (turingtest):

because 42/70>20/70